OpenStudy (anonymous):

You have 200 feet of fencing to enclose a rectangular plot that borders on a river. If you do not fence the side along the river, find the length and width of the plot that will maximize the area. What is the largest are that can be enclosed? A.) Find Vertex B.) Find dimesional L and W C.) Find maximum area

OpenStudy (saifoo.khan):

can u find it's example there?

OpenStudy (anonymous):

example 7 i think

OpenStudy (saifoo.khan):

Let;s begin.

OpenStudy (anonymous):

haha.. you've got the go ahead

OpenStudy (saifoo.khan):

A = LW 2W + L = 200 L = 200 - 2W A = W(-2W + 200) A = -2W^2 + 200W.

OpenStudy (anonymous):

thats it??

OpenStudy (saifoo.khan):

no... solving..

OpenStudy (anonymous):

ohh ok take ur time

OpenStudy (saifoo.khan):

i think im getting it, just a sec.

OpenStudy (anonymous):

okay

OpenStudy (saifoo.khan):

Let x be the length of the portion of the fence parallel to the river. Then the other two sides each have length (200 − x)/2, and the total area is....

OpenStudy (saifoo.khan):

we did a mistake in the last one.

OpenStudy (anonymous):

really? where

OpenStudy (saifoo.khan):

the vertex there was (12,0)

OpenStudy (saifoo.khan):

rest is correct.

OpenStudy (anonymous):

okay.. fixed it

OpenStudy (saifoo.khan):

here, vertex = (100,5000)

OpenStudy (saifoo.khan):

this is hard for a 16 year old. :O

OpenStudy (saifoo.khan):

the max area is 5000.

OpenStudy (saifoo.khan):

Mystery solved.. :D

OpenStudy (anonymous):

lol.. its hard for a 19 yr old

OpenStudy (anonymous):

wats B?

OpenStudy (saifoo.khan):

length = 100 width = 50.

OpenStudy (saifoo.khan):

that's B.

OpenStudy (anonymous):

how did u get 100,5000 for vertex. like the formula u used

OpenStudy (saifoo.khan):

-b/2a

OpenStudy (saifoo.khan):

i got 100. then insert that value in it again to get 5000

OpenStudy (saifoo.khan):

man! these were hard.

OpenStudy (anonymous):

wait.. A= -2 B=200 rite

OpenStudy (anonymous):

so i can be able to find the vertex

OpenStudy (saifoo.khan):

Nope. -b/2a -(-100)/2(0.5)

OpenStudy (saifoo.khan):

100

OpenStudy (anonymous):

can u put it like the way u had put the first one please

OpenStudy (anonymous):

if u don't mind